The square of Janet's age is 400 more than the square of the sum of Kim's and Sue's ages. Kims's and Sue's ages total 10 less than Janet's age. Find the square of the sum of the ages of Janet, Kim and Sue. Explain your reasoning.
Answer by edjones(8007)

Let x=sum of the ages of Kim and Sue, j=age of Janet.
j^2=x^2+400
j=x+10
.
(x+10)^2=x^2+400 Substitute x+10 fo j.
x^2+20x+100=x^2+400
20x=300
x=15
j=25
.
(25+15)^2
=40^2
=1600 the square of the sum of the ages of Janet, Kim and Sue.
.
Ed